OpenStudy (anonymous):

The volume of a hexagonal prism is 12square root 3 in3. What is the height of the prism if the base area is 3square root 3 in3? 4square root 3 in 36 in 36square root 3 in 4 in

OpenStudy (jdoe0001):

on the base area I gather your meant square units, that is \(\bf 3\sqrt{3}\ in^2\)

OpenStudy (jdoe0001):

anyhow, volume = length x height x width or for this prism will be base area x height \(\bf volume = 12\sqrt{3}\qquad \qquad \textrm{base area} = 3\sqrt{3}\ in^2\\ volume = \textit{base area} \times height\\ 12\sqrt{3} = 3\sqrt{3} \times height\) solve for "height"